21. Two 500 g blocks of wood are 2.0 m apart on a frictionless table. A 10 g bullet is fired at 400 m/s blocks. It passes all the way through the first block, then embeds itself in the second block. The speed of the first block immediately afterward is 6.0 m/s. What is the speed of the second block after the bullet stops in it?
Answer #1

The motion of the bullet and the blocks happen in the horizontal direction and we consider no external force acting along horizontal. If external force is zero, linear momentum of the system remains conserved. Here our system comprises of the two blocks and the bullet. We need to consider the intial and the final state of the system and equate the linear momentum.

Linear momentum, p = mv, where m is the mass of the object and v is the speed of that object.

Initial state:

10 g Bullet moving at a speed of 400 m/s while the two blocks are stationary. Hence, there will be momentum due to only the moving bullet. Hence,

pi = mbvb = 0.01 kg x 400 m/s = 4 kgm/s.........(1)

Final state:

Block 1 (m1 = 500 g = 0.5 kg) moves at v1 = 6 m/s. Bullet and Block 2 move together having collective mass,

m2 = 500g + 10 g = 510 g = 0.510 kg, move at say, v2. Hence,

pf = m1v1 + m2v2

pf = (0.5)x6 + (0.51)v2 = 3 + 0.51v2...........(2)

As momentum is conserved, equating (1) and (2)

4 = 3 + 0.51v2

0.51v2 = 1

Hence,

v2 = 1.96 m/s